how many no. lying between 10 and 1000 can be formed from the digits 1,2,3,4,5,6,7,8,9
Answer by josmiceli(19441)

They don't say if the numbers include
10 and 1000, but I'll assume they do.
1 through 9 is 9 numbers, and 0 is excluded
----------------------------------------
10, 20, 30, . . . 90 will be excluded ( 9 numbers )
------------------------------------------
10x and
1x0
are excluded where x = 0 - 9 ( 20 numbers )
------------------------------------------
20x and
2x0
are excluded ( 20 numbers )
----------------------------
Keep going, and this is 20 x 9 = 180 numbers
----------------------------
180 + 9 = 189 are excluded
------------------------------
10 through 1000 inclusive = 991 numbers
------------------------------
991 - 189 = 802
802 numbers can be formed from
1 through 9
------------------------------
Note that if 10 and 1000 are not included in the count
then 800 numbers can be formed
--------------------
feel free to get another opinion on this